please help me solve this proof
The formula learned for slope is m= y2-y1/x2-x1. use an algebraic proof to show that the slope can be claculated using the equation m= y1-y2/x1-x2

Answer by Alan3354(6059)

The formula learned for slope is m= y2-y1/x2-x1. use an algebraic proof to show that the slope can be claculated using the equation m= y1-y2/x1-x2
-------------
y1-y2 = -(y2-y1)
Same with the x terms
It's both NUM and DEN times -1, which is no change.